Solve the following equations:
10p = 100..............

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

We have to divide both sides of the equation by 10,
Then, we get,
= 10p/10 = 100/10
= p = 10

Thanks
Prachi Singh
2 years ago
Dear Student,
 
We can solve the following expression as follows :
10p = 100
P = 100/10
P = 10
 
Hope it helps!
Thank you